# Experimental Tests of Invariant Set Theory

### v2

@@ Abstract
We identify points of difference between Invariant Set Theory and standard quantum theory, and show that these lead to noticeable differences in predictions between the two theories. We design a number of experiments to test which of these predictions corresponds to our world. If these experiments were undertaken, they would allow us to investigate whether standard quantum theory or invariant set theory best describes reality. These tests can also be deployed on theories sharing similar properties (e.g. Penrose’s gravitational collapse theory).
@@ I Introduction
Invariant Set Theory (IST) attempts to unify these two disparate branches by using insight from Chaos Theory to create a fully local and determinate model of quantum phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. It does this by assuming that the universe is a determinate dynamical system evolving precisely on a fractal invariant set in state space. The natural metric to describe distances on a fractal set is the $p$-adic metric. This $p$-adic metric replaces the standard Euclidean metric of distance between states in state space. A consequence of this switch is that putative counterfactual states which lie in the fractal gaps of the invariant set are considered distant from states which do lie on the invariant set, even though from a Euclidean perspective such distances may appear small. Given the uncomputability of the possible states on any given fractal attractor, we cannot in advance distinguish states allowed and disallowed by this metric. Hence, in IST, quantum-scale phenomena appear random despite being deterministic.
$p$-adic numbers form a back-bone of modern number theory and as such provide a framework to describe quantum physics within a finite number-theoretic framework. An example of this is how IST provide an explanation for complementarity, a concept underpinning the uncertainty principle in quantum mechanics. In IST, complementarity is an emergent phenomenon arising from a number-theoretic property of trigonometric functions known as Niven’s theorem—that $\cos\phi$ is not a rational number when $\exp{i\phi}$ is a primitive $p$th root of unity. The complex Hilbert Space of standard quantum mechanics arises as a singular limit of invariant set theory when $p$ is set equal to infinity.
However, despite showing how key examples of quantum phenomena (like the sequential Stern-Gerlach effect, and Bell inequality violation) can be described deterministically, the theory deviates from standard quantum physics in some of its predictions—mainly in ways which stem from the $p$-adic metric being finite. In this paper, we give these key points of deviation, and investigate the extent to which these could be used to experimentally test the theory.
@@ II Invariant Set Theory
Palmer has not yet given a written dynamical law for IST. However, this is common for quantum foundational models, as we often only care about transition amplitudes between initial and final times. Further, these amplitudes are often between simple qubit observables (e.g., spin or polarisation states), meaning these models often don’t give or require a space-time evolution law either. Spekkens’s Toy Model is an example of such a model [20], having no dynamical evolution equation but still proving useful for considering foundational questions. Similarly to this model, IST in its current form exists to see what interesting insights we can get from considering incomputable counterfactual restrictions on state space.
@@
Obviously, thinking about rational numbers and restrictions on which states can mutually exist does not explain all results of quantum mechanics—for that, we would need other things, such as a dynamical law. However it does present a a mathematical mechanism for us being unable to simultaneously measure certain combinations of properties of quantum systems.
@@ III Entanglement Limits
In standard quantum theory, there is no limit to the number of quantum objects $n$ which can be $n$-partite maximally entangled (i.e., saturate the Coffman-Kundu-Wootters inequality for an $n$-partite state [21, 22]). However, in IST, there is. Here, we codify this limit, and design experiments using optics/noisy intermediate-scale quantum devices (NISQ devices) to probe it.
For this, we use the $M$-qubit W state [23, 24],
$\left|W^{M}\right\rangle=\frac{1}{\sqrt{M}}\sum^{M-1}_{i=0}\left|0\right\rangle^{\otimes i}\left|1\right\rangle\left|0\right\rangle^{\otimes(M-1-i)}$ (3)
(where $\left|\psi\right\rangle^{\otimes M}$ is the tensor product of $\left|\psi\right\rangle$ with itself $M$ times). For instance, the W state where $M=3$ is
@@
The W state is a maximally entangled state of $M$ qubits—and in standard quantum theory, there is no limit to how high $M$ can be.
However, in IST, the finiteness of the $p$-adic metric provides a limit to the number of qudits that can be maximally entangled. For multiple-qubit entanglement, this limit is codified in [11] as a maximum of $\log_{2}N$ qubits being able to be maximally entangled, in a $p$-adic system where the equatorial great circle of the Bloch sphere consists of $N$ equally-spaced discrete points.
[…]
A system of maximally-entangled photon-vacuum qubits can be created using a single photon and a number of mirrors and 50:50 beamsplitters, as shown in Fig.1. This naturally forms a W state across $M$ qubits, and, by standard quantum theory, we should potentially be able to extend this set-up to $M\rightarrow\infty$. However, this disagrees with IST, which limits to a maximum of $M=\log_{2}N$ entangled qubits, where the two orthogonal spherical dimensions of the Bloch sphere ($\theta$ and $\phi$) are each discrete in $N$ divisions. While $p$ is expected to be very large, each qubit will only have been affected by $I=\log_{2}\log_{2}N$ beamsplitters, so, for realistic experimental beamsplitter loss of $0.1\%$, the chance of losing a given qubit to decoherence only reaches $1\%$ once the system has entangled over 1000 qubits, which is only possible by IST if $N\geq 10^{250}$ (as we show in Fig.2). Further, an advantage of the W state is, even if decoherence effectively measures one of the qubits, so long as the result is 0 (the photon isn’t in that mode), this collapse leaves the remaining qubits still maximally entangled in the $(M-1)^{th}$ W state.
Figure 2: The survival probability of each qubit for a given set of entangled qubits created using the experiment in Fig. 1, and the maximum number of entangled qubits that can be created in a version of IST where the $p$-adicity causes the Bloch sphere to be split into $N$ divisions in each angular direction. This shows how this beamsplitter experiment allows us to test this entanglement limit for very high-$p$ versions of IST, due to the comparative lack of loss-induced decoherence on the W state created.
@@
In the same way, for the $I^{th}$ iterate, consisting of $M=2^{I}$ qubits, using linear optical components one can project the eventual state into one of the $2^{I}$ phase permutations of $\left|W^{M}\right\rangle$, and so detect with certainty that a pure entangled state of $M$ qubits was created. Interestingly, preparing these states to certify entanglement requires each optical mode to again only interact with $I$ beamsplitters, to allow us to certify $M=2^{I}$-qubit entanglement, which simply squares the survival probability. This means for 1000 qubits, it becomes ${}98\%$ rather than ${}99\%$. Given the resilience of the overall state to loss-induced decoherence, and the fact that Lougovski et al show this certification method also allows us to detect any entangled states of fewer than $M$ qubits, this loss probability poses very little issue to our test of IST. Further, despite the loss, the total number of surviving (maximally-entangled) qubits tends to infinity as $I$ tends to infinity, rather than peaking at a certain value.
@@
Figure 3: Type I spontaneous parametric down-conversion (SPDC) source for the generation of pairs of position-and-momentum-entangled photons, as given by Rarity and Tapster [29]. The generated position of each photon on the cone can be viewed as a W state of arbitrary number of qubits $N$, and so the system of the two photons is a double-W state of $2N$ qubits. This arbitrary number of qubits $N$ can be lower-bounded as the resolution of a circular single-photon position detector array used to detect where on the circle each photon is emitted.
@@
This W state-based experimental analysis of IST can be extended by looking at an experiment such as Rarity and Tapster’s, where a pair of photons are generated in a cone of possible positions. Here, the angular position of one photon is anti-correlated with the position of the other [29]. We show this in Fig. 3. Considering just one photon in the cone, this is equivalent to a W state where $M$ is the number of sectors into which you subdivide the cone. Adding a second photon, position-entangled with the first, doubles the number of entangled qubits in the system.
@@
Rarity and Tapster also give a way to prove these photons are entangled: by interfering them to violate a Bell inequality. However, as this is done assuming their position is a continuous variable, we need to adjust it to prove just how large it holds for as discrete variables.
@@
This can be done by making a set of $2M$ apertures on the circumference of the cone, and splitting the ring into two half-circumferences. After this, similarly to what we do in Fig. 1, we can iteratively combine adjacent apertures to get position-momentum entanglement between adjacent apertures. Once this projects to equal superpositions across all $M$ apertures on each half-circle, we can record detected position for each half-circle’s photon. By comparing the final detected position between upper half-circumference and lower half-circumference, and seeing if they still correlate, we can confirm this double-WM state.
@@
IV No Continuous Variables
A second, related implication of IST is that it permits no continuous quantum variables. As the $p$-adic metric used in IST is necessarily finite-dimensional, the space of states allowed must also be finite. Since we can lower bound the number of states allowed as the dimension of the Hilbert space we use (to replicate classical information theory), we can say that, the existence of a qudit of dimension $d$ implies a state space of at least dimension $d$ (e.g. a qubit requires at least two distinct states: 0 and 1; a qutrit requires 3 states: 0, 1 and 2, etc…). Hardy extends this argument, saying that, to satisfy his axioms for quantum theory, between any two pure states in a system, there needs to be a continuous reversible transformation available on a system that goes from one to the other. To allow this, Hardy argues a qudit of dimension $d$ requires a state space of dimension $d^{2}$ [30].
In standard quantum physics, many variables are continuous (e.g., position, momentum, electric field strength, and time) [32]. Therefore, for IST to hold true, all of these variables would actually need to be discrete: of finite (but very high) dimension. While some models hold one or another of these variables to be continuous (e.g. space-time in Loop Quantum Gravity [33] and certain toy models of the Universe [34]), the idea that all ‘continuous’ variables are actually discrete would be controversial.

### v3

@@ Abstract
We identify points of difference between Invariant Set Theory and standard quantum theory, and show that these lead to noticeable differences in predictions between the two theories. We design a number of experiments to test which of these predictions corresponds to our world. If these experiments were undertaken, they would allow us to investigate whether standard quantum theory or invariant set theory best describes reality. These tests can also be deployed on theories sharing similar properties (e.g., Penrose’s gravitational collapse theory).
@@ I Introduction
Invariant Set Theory (IST) attempts to unify these two disparate branches by using insight from Chaos Theory to create a fully local and determinate model of quantum phenomena [1, 2, 3, 4]. It does this by assuming that the universe is a determinate dynamical system evolving precisely on a fractal invariant set in state space. The natural metric to describe distances on a fractal set is the $p$-adic metric (a fractal metric, with different properties to the Euclidean metric with which define spatial distance).
In IST, this $p$-adic metric replaces the standard Euclidean metric of distance between states in state space. A consequence of this switch is that putative counterfactual states which lie in the fractal gaps of the invariant set are considered distant from states which do lie on the invariant set, even though from a Euclidean perspective such distances may appear small. In IST, this is used as an argument for these states being counterfactually restricted [5]—in the same way as spaces between points on a “snap-to” grid on a computer are inaccessible, these $p$-adically distant states are inaccessible in IST. Given which points are on any given fractal attractor is formally uncomputable, we cannot in advance distinguish states allowed and disallowed by this metric. Hence, in IST, quantum-scale phenomena appear random despite being deterministic.
$p$-adic numbers form a back-bone of modern number theory [6] and as such provide way for us to use sophisticated tools from finite number theory to describe quantum physics. An example of this is how IST explains complementarity, a concept underpinning the uncertainty principle in quantum mechanics. In IST, complementarity is an emergent phenomenon arising from Niven’s theorem, a number-theoretic property of trigonometric functions. The theorem states that $\cos\phi$ is not a rational number when $\exp{i\phi}$ is a primitive $p$th root of unity (i.e., when $\phi$ is not 0 or some integer multiple of $\pi$). Notably, the complex Hilbert Space of standard quantum mechanics seems to arise as a singular limit of invariant set theory when $p$ is set equal to infinity.
However, despite showing how key examples of quantum phenomena (like the sequential Stern-Gerlach effect, and Bell inequality violation, in [4]) can be described deterministically, the theory deviates from standard quantum physics in some of its predictions—mainly in ways which stem from the $p$-adic metric being finite. While there have been some attempts to challenge IST mathematically [7], these empirical differences have not yet been properly considered. In this paper, we give these key points of deviation, and investigate the extent to which these could be used to experimentally test the theory.
@@ II Invariant Set Theory
There is not yet a written dynamical law for IST. However, this is common for quantum foundational models, as we often only care about transition amplitudes between initial and final times. Further, these amplitudes are often between simple qubit observables (e.g., spin or polarisation states), meaning these models often don’t give or require a space-time evolution law either. Spekkens’s Toy Model is an example of such a model [14], having no dynamical evolution equation but still proving useful for considering foundational questions. Similarly to this model, IST in its current form exists to see what interesting insights we can get from considering incomputable counterfactual restrictions on state space.
@@
Obviously, thinking about rational numbers and restrictions on which states can mutually exist does not explain all results of quantum mechanics—for that, we would need other things, such as a dynamical law. However it does present a mathematical mechanism for us being unable to simultaneously measure certain combinations of properties of quantum systems.
@@
Instead, we use the fact that we can build a representation of a qubit obeying the discretisation conditions given in Eq. 2 using a bit-string of length $p$. To explore empirical differences between IST and standard quantum mechanics, we following [4] to do so. Note though that we only describe the elements of this which are necessary to motivate the experimental differences we discuss below—if you are more interested in how such a bit-string model of qubits in IST can be used to demonstrate complementarity, or to violate a Bell inequality, we strongly suggest you read through [4]. In this bit-string, amplitude (treated as rotation from the top pole of the Bloch sphere in our basis of choice) can be represented by proportion of bits taking one value; and phase (treated as rotation around the circumference of the Bloch sphere) can be represented as cyclic permutation of this bit-string from the scenario where it is, in-order, first all 0-bits, then all 1-bits—i.e., if the value of the $i^{t}h$ bit in the bit string is $a_{i}$, phase $\phi=2\pi n/p$ causes cyclic permutation:
$\zeta^{n=\phi p/2\pi}\{a_{1},a_{2},...a_{p}\}=\{a_{n+1},a_{n+2},...a_{p},a_{1},a_{2},...a_{n}\}$ (3)
This makes sense, given, at either pole of the Bloch sphere (relative to this representation), the bit-string is either all 0s or all 1s, so a cyclic permutation of the bits in the string does nothing to the bit-string. The various unitary operations which can be done on a qubit can all be represented as rotations on the Bloch sphere, so can all be represented as changes to the values of bits in the bit-string (increasing or decreasing the relative proportion of 0s and 1s to change the latitude of the state on the Bloch sphere), to the cyclic permutation of the position of the bits in the string relative to the “phaseless” bit-string
$\{0_{1},0_{2},...0_{p-m},1_{p-m+1},...1_{p}\}$ (4)
where, for an arbitrary qubit $\left|\psi\right\rangle=\alpha\left|0\right\rangle+\beta\left|1\right\rangle$, $|\beta|^{2}=m/p$.
This shows how IST uses bit-strings to represent a single qubit, but what about multiple, potentially-entangled qubits?
To do this, for $J$ qubits, IST says we need to take the Cartesian product of $J$ bit-strings, each of length $p$. Where the “phaseless” form of the first bit-string is partitioned as $p-m_{1}$ 0s, then $m_{1}$ 1s, the “phaseless” form of the second bit-string is here partitioned $p-m_{2}$ 0s, then $m_{2}-m_{1}$ 1s, then $m_{3}$ 0s, then $m_{1}-m_{3}$ 1s. This means the first sub-string of 0s then 1s of the second bit-string is the same length as the chain of 0s of the first bit-string ($p-m_{1}$), and the second sub-string of the second bit-string is the same length as the chain of 1s in the first bit-string ($m_{1})$. When there is exactly the same ratio of 0s to 1s in the first sub-string as there is in the second sub-string, then obviously the state of the preceding string (or rather, preceding qubit) has no effect on the state of that qubit—therefore, the two qubits are unentangled. However, when the state is maximally affected (as in, there are say no 0s in the first sub-string, and no 1s in the second sub-string), then this is effectively a maximal entanglement. Beyond this phaseless case, we also now have three different cyclic permutation operations—one on the first-bit-string, as before, but also one on the first sub-string of the second bit-string, and one on the second sub-string of the second bit-string. As well as in correlation between weightings of 0s and 1s, entanglement can also manifest in correlations imposed by these cyclic permutations (i.e., there being difference between the number of times the cyclic permutation operator is applied to the first sub-string of the second bit-string, and the number of times it is applied to the second sub-string). For the two qubits, this gives us 6 free integer parameters. Given quantum mechanics, for two qubits, requires a state space $\mathcal{S}^{6}$, this should (up to a certain resolution) allow us to represent two entangled qubits. Note though that, due to the partitioning of the second qubit into two sub-strings, we effectively half the resolution which we had previously for a single qubit.
This subdivision (in the phaseless case) goes similarly for a third bit-string (with has four sub-strings, two for each of the two sub-strings of the second bit-string), and for the fourth bit-string (which has 8 sub-strings, two for each of the four sub-strings of the second bit-string), and so on, for as many qubits as we wish to add to the system. Entanglement between any two of the qubits can manifest through differences in relative weighting of 0s and 1s in certain sub-strings, differences in number of times cyclic permutation is applied in certain substrings, or a combination of these two effects. However, the more qubits are added to the system, the lower the resolution with which the systems are able to be represented within the $p$-length bit-strings, and so the smaller the amount of entanglement-induced correlation can be represented on the system. In order for us to be able to use Niven’s Theorem to reproduce complementarity for a single qubit, [4] shows that we need a bit string of length $p=2^{M}$, where $M\geq 2$. Similarly, given the progressive partitioning of bit-strings into sub-strings as we add potentially-correlated qubits to the system, for our system of $J$ qubits to behave in any way quantum-mechanically together, we need $M\geq 2J$, and so $p\geq 2^{2J}$. Put another way, IST presents a limit of $J=\log_{2}(p)/2$ qubits which can be entangled together.
@@ III Entanglement Limits
[…]
In standard quantum theory, there is no limit to the number of quantum objects $n$ which can be $n$-partite maximally entangled (i.e., saturate the Coffman-Kundu-Wootters inequality for an $n$-partite state [15, 16]). However, in IST, there is (as we saw above). Here, we codify this limit, and design experiments using optics/noisy intermediate-scale quantum devices (NISQ devices) to probe it.
For this, we use the $J$-qubit W state [17, 18],
$\left|W^{J}\right\rangle=\frac{1}{\sqrt{J}}\sum^{J-1}_{i=0}\left|0\right\rangle^{\otimes i}\left|1\right\rangle\left|0\right\rangle^{\otimes(J-1-i)}$ (5)
(where $\left|\psi\right\rangle^{\otimes J}$ is the tensor product of $\left|\psi\right\rangle$ with itself $J$ times). For instance, the W state where $J=3$ is
@@
The W state is a maximally entangled state of $J$ qubits—and in standard quantum theory, there is no limit to how high $J$ can be.
However, in IST, the finiteness of the $p$-adic metric provides a limit to the number of qudits that can be maximally entangled. As mentioned above, for multiple-qubit entanglement, this limit is codified in [4] as a maximum of $\log_{2}p$ qubits being able to be maximally entangled, in a $p$-adic system where the equatorial great circle of the Bloch sphere consists of $p$ equally-spaced discrete points.
A system of maximally-entangled photon-vacuum qubits can be created using a single photon and a number of mirrors and 50:50 beamsplitters, as shown in Fig. 1. This naturally forms a W state across $J$ qubits, and, by standard quantum theory, we should potentially be able to extend this set-up to $J\rightarrow\infty$. However, this disagrees with IST, which limits to a maximum of $J=\log_{2}J$ entangled qubits, where the two orthogonal spherical dimensions of the Bloch sphere ($\theta$ and $\phi$) are each discrete in $J$ divisions. While $p$ is expected to be very large, each qubit will only have been affected by $I=\log_{2}\log_{2}J$ beamsplitters, so, for realistic experimental beamsplitter loss of $0.1\%$, the chance of losing a given qubit to decoherence only reaches $1\%$ once the system has entangled over 1000 qubits, which is only possible by IST if $J\geq 10^{250}$ (as we show in Fig. 2). Further, an advantage of the W state is, even if decoherence effectively measures one of the qubits, so long as the result is 0 (the photon isn’t in that mode), this collapse leaves the remaining qubits still maximally entangled in the $(J-1)^{th}$ W state.
Figure 2: The survival probability of each qubit for a given set of entangled qubits created using the experiment in Fig. 1, and the maximum number of entangled qubits that can be created in a version of IST where the $p$-adicity causes the Bloch sphere to be split into $J$ divisions in each angular direction. This shows how this beamsplitter experiment allows us to test this entanglement limit for very high-$p$ versions of IST, due to the comparative lack of loss-induced decoherence on the W state created.
@@
In the same way, for the $I^{th}$ iterate, consisting of $J=2^{I}$ qubits, using linear optical components one can project the eventual state into one of the $2^{I}$ phase permutations of $\left|W^{J}\right\rangle$, and so detect with certainty that a pure entangled state of $J$ qubits was created. Interestingly, preparing these states to certify entanglement requires each optical mode to again only interact with $I$ beamsplitters, to allow us to certify $J=2^{I}$-qubit entanglement, which simply squares the survival probability. This means for 1000 qubits, it becomes ${}98\%$ rather than ${}99\%$. Given the resilience of the overall state to loss-induced decoherence, and the fact that Lougovski et al show this certification method also allows us to detect any entangled states of fewer than $J$ qubits, this loss probability poses very little issue to our test of IST. Further, despite the loss, the total number of surviving (maximally-entangled) qubits tends to infinity as $I$ tends to infinity, rather than peaking at a certain value.
IV No Continuous Variables
A second, related implication of IST is that it permits no continuous quantum variables. As the $p$-adic metric used in IST is necessarily finite-dimensional, the space of states allowed must also be finite. This can be observed in the bit-string model above through the length limit $p$ per bit-string, and the binary value of each bit—similarly to the finite resolution this allows for the representation of the qubit state on the Bloch sphere, this would only allow a finite coarse-grained value for any continuous variable, rather than an exact specification of a value on a continuum. Since we can lower bound the number of states allowed as the dimension of the Hilbert space we use (to replicate classical information theory), we can say that, the existence of a qudit of dimension $d$ implies a state space of at least dimension $d$ (e.g. a qubit requires at least two distinct states: 0 and 1; a qutrit requires 3 states: 0, 1 and 2, etc…). Hardy extends this argument, saying that, to satisfy his axioms for quantum theory, between any two pure states in a system, there needs to be a continuous reversible transformation available on a system that goes from one to the other. To allow this, Hardy argues a qudit of dimension $d$ requires a state space of dimension $d^{2}$ [23].
In standard quantum physics, many variables are continuous (e.g., position, momentum, electric field strength, and time) [25]. Therefore, for IST to hold true, all of these variables would actually need to be discrete: of finite (but very high) dimension. While some models hold one or another of these variables to be continuous (e.g. space-time in Loop Quantum Gravity [26, 27, 28] and certain toy models of the Universe [29]), the idea that all ‘continuous’ variables are actually discrete would be controversial.
@@ IV No Continuous Variables
Figure 3: Type I spontaneous parametric down-conversion (SPDC) source for the generation of pairs of position-and-momentum-entangled photons, as given by Rarity and Tapster [30]. The generated position of each photon on the cone can be viewed as a W state of arbitrary number of qubits $J$, and so the system of the two photons is a double-W state of $2J$ qubits. This arbitrary number of qubits $J$ can be lower-bounded as the resolution of a circular single-photon position detector array used to detect where on the circle each photon is emitted.
@@
Probing this difference between discrete and continuous treatments of certain variables presents yet another way to test IST—seeing for just how fine-grained a discretisation the experimental effects we would expect (from entangled continuous variables in standard quantum mechanics) still manifest when we instead treat the system as fundamentally discrete and modelled through IST, then probing the system at the level where these predictions diverge. The W state-based experimental analysis we give above can be extended in this way by looking at an experiment such as Rarity and Tapster’s, where a pair of photons are generated in what is assumed in standard quantum mechanics to be a continuous cone of possible positions. Here, the angular position of one photon is anti-correlated with the position of the other [30]. We show this in Fig. 3. Considering just one photon in the cone, this is equivalent to a W state where $J$ is the number of sectors into which you subdivide the cone. Adding a second photon, position-entangled with the first, doubles the number of entangled qubits in the system.
@@
Rarity and Tapster also give a way to prove these photons are entangled: by interfering them to violate a Bell inequality. However, as this is done assuming their position is a continuous variable, we need to adjust it to identify just what dimension of discrete variable this seemingly-maximal entanglement provably holds for.
@@
This can be done by making a set of $2J$ apertures on the circumference of the cone, and splitting the ring into two half-circumferences. After this, similarly to what we do in Fig. 1, we can iteratively combine adjacent apertures to get position-momentum entanglement between adjacent apertures. Once this projects to equal superpositions across all $J$ apertures on each half-circle, we can record detected position for each half-circle’s photon. By comparing the final detected position between upper half-circumference and lower half-circumference, and seeing if they still correlate, we can confirm this double-WJ state.
@@
According to IST, given the finite resolution available for the representation of both the states of the two photons and the correlations between them, eventually, at a fine enough resolution, we should begin to see the entanglement-induced correlations between the two photons breaking down. As we move to the number of apertures $J$ being of the same order as $p$, IST predicts all entanglement between the two photons should have degraded. However, standard quantum mechanics obviously predicts no aperture-number-dependent degradation of the entanglement-induced correlation. This therefore presents another test we can use to distinguish IST and standard quantum mechanics.